# D-Gril: End-to-End Topological Learning with $2$-parameter Persistence

### v2

@@ 5.3 Benchmark Graph Datasets
D-Gril can be used more generally for filtration learning on graph datasets. We perform a series of experiments with benchmark graph datasets such as Mutag, Proteins, Dhfr, Cox2 Morris et al. (2020) and compare with existing multiparameter persistence methods. Details about the datasets are given in Appendix B. We report the results in Table 3. For a valid comparison, we use a $3$-layer MLP as the classifier for all the multiparameter signatures. Hence, the numbers reported look different from the ones in Xin et al. (2023), Carrière & Blumberg (2020) where the authors consider XGBoost (Chen & Guestrin, 2016). In fact, in Xin et al. (2023), the authors show that XGBoost performs much better than a $3$-layer MLP classifier for Gril vectors on these datasets. We can see from the table that learning the bifiltration function seems to perform better than multiparameter persistence methods on popular choices of bifiltration functions on most datasets. We can also see that D-Gril performs better than Gril, supporting our argument for an end-to-end learning framework. We report the training times in Table 4. We can see from the table that the approach is feasible and can be used in practice.
Dataset Time(hh:mm:ss)
MUTAG 00:05:08
COX2 00:11:34
DHFR 00:16:48
PROTEINS 00:30:03
Table 4: Reported times are training times per fold averaged over 5 folds that we used for training. All the experiments have been performed on a machine with AMD EPYC 7313 16-Core Processor and NVIDIA A10 GPU.
In order to visualize the learnt bifiltration function, we plot it and the D-Gril vectors for $H_{0}$ and $H_{1}$ and compare it with the plots of Heat-Kernel Signature-Ricci Curvature (HKS-RC) bifiltration function and corresponding Gril vectors on two randomly selected graph instances from the Proteins dataset. The figures are shown in Appendix B.
@@ 6 Conclusion
In this paper, we propose a differentiable framework using Gril, a $2$-parameter persistent homology based vectorization. We show that stochastic sub-gradient descent converges almost surely on Gril, and we compute an explicit formula for the differential of Gril. We use this formula to back-propagate gradients through the Gril computation, yielding the differentiable framework D-Gril. We use D-Gril to show that adding topological information in an end-to-end manner is beneficial for bio-activity prediction. Further, we show that D-Gril can be used in a more general setting, for filtration learning on graphs. Our results indicate that learnt bifiltration functions on graphs, generally, give a better performance than the standard choices for bifiltration functions. This indicates, and we believe, that learning filtration functions can prove to be more informative for topological methods in machine learning. We hope that this work sparks interest and motivates research in this direction.

### v3

@@ 5.3 Benchmark Graph Datasets
D-Gril can be used more generally for filtration learning on graph datasets. We perform a series of experiments with benchmark graph datasets such as Mutag, Proteins, Dhfr, Cox2 Morris et al. (2020) and compare with existing multiparameter persistence methods. Details about the datasets are given in Appendix B. We report the results in Table 3. For a valid comparison, we use a $3$-layer MLP as the classifier for all the multiparameter signatures. Hence, the numbers reported look different from the ones in Xin et al. (2023), Carrière & Blumberg (2020) where the authors consider XGBoost (Chen & Guestrin, 2016). In fact, in Xin et al. (2023), the authors show that XGBoost performs much better than a $3$-layer MLP classifier for Gril vectors on these datasets. We can see from the table that learning the bifiltration function seems to perform better than multiparameter persistence methods on popular choices of bifiltration functions on most datasets. We can also see that D-Gril performs better than Gril, supporting our argument for an end-to-end learning framework. We also see that D-Gril performs better than Mpsm. The training times for the two end-to-end learning methods, D-Gril and Mpsm are reported in Table 4. We can see that the training times for D-Gril are improved by about $40\%$ compared to the ones of Mpsm. We refer the reader to Figure  4 for visualizations of bifiltration functions learnt by D-Gril.
@@ 6 Conclusion
In this paper, we propose a differentiable framework using Gril. We show that stochastic sub-gradient descent converges almost surely on Gril, and we compute an explicit formula for the differential of Gril. We use this formula to back-propagate gradients through the Gril computation, yielding the differentiable framework D-Gril. We use D-Gril to show that adding topological information in an end-to-end manner is beneficial for bio-activity prediction. Further, we show that D-Gril can be used in a more general setting, for filtration learning on graphs. Moreover, we show that D-Gril is faster and lighter, both in theory and in practice, than Mpsm. Our results indicate that learnt bifiltration functions on graphs, generally, give a better performance than the standard choices for bifiltration functions. This indicates, and we believe, that learning filtration functions can prove to be more informative for topological methods in machine learning. We hope that this work sparks interest and motivates research in this direction.